# NarraBench: A Comprehensive Framework for Narrative Benchmarking

### v1

@@ 1 Introduction
Narratives are omnipresent in daily life. They can be used to entertain, inform, persuade, and maintain shared beliefs in communities across generations. Designing large language models (LLMs) that can both understand and generate narrative communication is of utmost importance to building artificially intelligent systems useful for human beings.
Figure 1: The twelve primary narrative features of the NarraBench taxonomy coloured according to the Big-4 narrative dimensions (story, narration, discourse, and situatedness). Cumulative goodness-of-fit values of existing benchmarks as determined by our benchmark survey are included as shading.
Recent years have seen a wave of new benchmarks that use narratives to test long context understanding and models’ theory of mind (Kim et al., 2023; Karpinska et al., 2024; Chen et al., 2024b; Wang et al., 2025a). Many standard LLM benchmarks also draw on narrative skills, even when they are not presented as such: “Needle in a haystack” tests show models can correctly discern what is relevant to the present contextual frame (Kamradt, 2025), and summarization tasks require models to integrate information across long narrative contexts to identify important through-lines (Lin, 2004).
Several large-scale benchmarks contain a significant amount of narrative content, regardless of their original design focus. For example, we estimate approximately 1 in 7 BIG-bench questions contains a story (Srivastava et al., 2023), 1 in 4 in HellaSwag (Zellers et al., 2019), and 1 in 3 in MMLU (Hendrycks et al., 2021). [^1] All datasets were classified on a per-document basis with the BERT-based classifier from Antoniak et al. (2024). Correctly answering these questions means language models must be capable of some form of narrative understanding.
Nevertheless, benchmarks that explicitly and systematically test for narrative understanding are rare and often lack theoretical consistency in their principal goals and methodological foundations. For example, existing benchmarks are nearly always oriented around single-answer frameworks, overlooking the way many narrative understanding tasks are fundamentally interpretive (i.e., perspectival) rather than deterministic in nature. Similarly, benchmarks tend to focus overwhelmingly on story-level content, missing more complex dimensions of narrative communication that include discourse structure and narrative perspective.
@@ 2 Related Work
Early computational work modelled narratives as stochastic processes (Kahn, 1973) or scripts known as Fillmorean frames (Schank and Abelson, 1975) containing chronological chains of events composed of entities and their actions (Chambers and Jurafsky, 2008, 2009, 2010; Balasubramanian et al., 2013; Reiter et al., 2014). Advancements in NLP led to increasingly sophisticated narrative extraction methods that made use of topic models, manually derived features, and associated rule mining (Antoniak et al., 2019; Belyy and Van Durme, 2020; Lyu et al., 2021; Zhang et al., 2023). This approach was extended to plot arc extraction, newly reinterpreted as sentiment over time (Hogan, 2011; Reagan et al., 2016; Somasundaran et al., 2020; Elkins, 2022; Knight et al., 2024), and social network analysis, where entities and their actions were rendered as graphs (Roemmele, 2019; Sims and Bamman, 2020; Tangherlini et al., 2020). Further work has focused on higher-level understanding including narrativity detection Antoniak et al. (2025); Piper and Bagga (2025) and narrative intent understanding Zhu et al. (2023). As LLMs are increasingly used to assess and generate narratives, now is an optimal time to codify the diverse array of tasks that fall under the heading of narrative understanding.
@@
Benchmarks are the principal infrastructure of empirical NLP, enabling standardized comparison and shared progress measures across models. Recent years have seen an explosion of benchmarks designed to evaluate long-context understanding (Chang et al., 2024; Karpinska et al., 2024; Wang et al., 2025a; Thai and Iyyer, 2025) and summarization performance (Fabbri et al., 2021; Zhao et al., 2022; Subbiah et al., 2025). A smaller set of benchmarks focus on explicitly narrative phenomena—including dialogue (Vishnubhotla et al., 2022), plot arcs (Chun, 2021), or genre classification (Worsham and Kalita, )—but general frameworks for narrative understanding remain scarce. Evaluation methods also remain limited; most rely on multiple-choice questions for efficiency, though MCQs are increasingly recognized as unreliable indicators of model comprehension (Sap et al., 2019; Wang et al., 2024; Balepur et al., 2024; Alzahrani et al., 2024; Wang et al., 2025b). Recent alternatives such as LLM-as-a-Judge evaluation (Li et al., 2024; Tan et al., 2025; Wang et al., 2025c) and distributional or multi-annotator scoring (Ni et al., 2024; Meister et al., 2024) point toward richer, less rigid evaluation pipelines, but have not yet been applied to narrative domains.
The NarraBench framework builds on these emerging perspectives by systematizing the conceptual space of narrative understanding and identifying the missing regions and evaluation criteria that current benchmarks overlook.
@@ 3 Theoretical Foundations
Narrative theory has developed, over more than a century, increasingly sophisticated qualitative descriptions of narratives across different mediums and contextual settings. We draw on this resource to build our taxonomy.
@@ 3.1 The Elements of Narrative
Our base definition of narrative communication is drawn from Piper et al. (2021). At its most elementary level, a narrative can be said to occur when all of the following criteria are met:
@@ A Someone
Narrative theory has additionally foregrounded the centrality of change or conflict as an elementary component of narrative communication Herman (2009), through concepts like “change of state” Prince (2012), “canonicity/breach” Bruner (1991), denouement Freytag (1895), or Aristotle’s foundational emphasis in the Poetics (c. 330 bce) on beginning, middle, and end structure. At a root level, narratives foreground temporal difference with respect to agent-centred events Genette (1980); Sternberg (1992); Ricoeur (2012).
@@ 3.2 Higher-Order Narrative Dimensions
Figure 2: NarraBench’s primary theoretical foundations from Genette (1980) and Herman (2009).
To integrate these elements into a broader framework, we combine Genette’s narrative triangle with Herman’s notion of situatedness (Figure 2). The Russian formalists Tomashevsky (1965) were the first to distinguish between story (“what happened”) and discourse (“how it was told”), to which Genette (1980) added a third dimension called narration, which refers to the perspectival dimensions of narrative communication produced by the narrator’s voice (“who speaks”).
Genette then introduced three further terms to capture the relationship between these dimensions, tense, mood, and voice. Genette extrapolates from the linguistic meanings of these terms to capture specific narratological features. These include aspects of time and the ordering of events (tense); the relationship between eventfulness and description (mood); and aspects related to perspective, such as point of view, dialogue, and focalization (voice). Herman (2009) importantly adds a fourth dimension, situatedness, which captures the social dimensions of narrative (medium, social context, interactive dimensions). We combine these elementary and higher-level dimensions to structure our taxonomy described in the next section.
@@ 4 The NarraBench Taxonomy
The second level of our taxonomy focuses on evaluation criteria. These include textual scale (local, global, meso), mode (discrete, progressive, holistic judgments), and the expected variance of potential answers (deterministic, consensus, perspectival). We present the taxonomy in its entirety in Table 2.
@@ 4.1.1 STORY
Story captures all aspects of a narrative that relate to storyworld content. Within this dimension we include six features:
@@
Narratives foreground the lived experiences of agentic entities (i.e., characters) Fludernik (2002). Agent detection and inventorying character names are an essential first step. Higher-level synthesis of characters around their narrative roles and attributes is necessary for understanding the functional position of characters within narratives Propp (1968). We also include theory-of-mind related tasks such as narrative emotion labelling and character motivation understanding. Interpreting agent intentionality and inner experience is a central component of narrative understanding Zunshine (2006); Mar (2011); Oatley (2016).
@@
Agents in a narrative exist in a particular social configuration. Here we divide social networks into three core aspects: interactions (local social events), connections (who does each character know), and relationships (the social types of connections, including employment, family, civil, romantic, and friendship-based connections) Labatut and Bost (2019).
@@
Narratives are structured around agents who experience change over time. Such change arises through action, where an action and its corresponding effect together form a causal frame. While narratologists debate the precise boundaries of a frame within a story, it is broadly accepted that reconstructing the underlying chain of actions and their causal relations is central to recovering the narrative’s event structure or schemas Schank and Abelson (1975); Chambers and Jurafsky (2009).
@@
According to Kukkonen (2014), “The term plot designates the ways in which the events and characters’ actions in a story are arranged and how this arrangement in turn facilitates identification of their motivations and consequences.” In a technical sense, plot sits between story and discourse, but for our purposes we consider it as part of the story dimension. Plot can thus be subdivided into topics (core thematic concerns), plot lines (sub-plots), key structural forces like narrative obstacles or conflicts, central value lessons (i.e., morals), as well as larger schematic structures like archetypes.
@@
Kukkonen (2014) makes a core distinction between plot as a function of “global structure” (our primary definition above) and “progressive structuration,” that is, aspects of plot that depend on the sequencing of events for their meaning. We reserve these progressive aspects for the discourse dimension, with the exception of plot arcs, which capture structural elements related to events such as fortune Reagan et al. (2016); Elkins (2022); McAdams et al. (2001), reversal Knight et al. (2024), and denouement Freytag (1895). Understanding the reliance or inability of LLMs to produce certain plot arc structures is important for shaping reader expectations.
@@
The spatial distribution of perception and movement within narratives is a key aspect of defining audiences’ cognitive orientation. Where narrative theory has introduced concepts such as the “chronotope” Bakhtin (2010) or “possible worlds” Ryan et al. (2016), we use the more integrative term of setting to capture both the holistic understanding of narrative space and the particular location of a given narrative moment.
@@ 4.1.2 NARRATION
Narratives require narrators. According to Genette (1980), “who speaks” plays a central role in governing the selection and communication of narrative content, which will in turn impact how audiences interpret the narrative.
@@
We use style to capture rhetorical and intertextual aspects of narratives that bear on their meaning. Allusion refers to inter-textual references Kristeva (1980); Genette (1997), while figurative language and imageability are concerned with the symbolic and/or perceptual dimensions of narrative communication. We also include complexity as a measure of reading level and social capital and evaluative language as a way of capturing narrative meta-awareness Labov and Waletzky (1997).
@@ 4.1.3 DISCOURSE
Narratives take place over time. But time is not necessarily represented in a chronological manner, nor is time represented in an equitable manner. Narrators may decide to dedicate greater fidelity to certain events, or reflect back on prior events or foreshadow future events. Following Genette (1980), we focus on both duration and order with respect to time.
@@
How information is revealed to the reader over the course of narrative time fundamentally shapes readers’ affective state Brewer and Lichtenstein (1982). Here we focus on the classic tripartite scheme of reader suspense, curiosity, and surprise to capture the assessment of narrative revelation over time.
@@ 4.2 Evaluation Criteria
This dimension captures assumptions about interpretive freedom. Some features may require single, correct answers, such as a character name or a scene location. We label these deterministic. Other features are more open-ended, such as narrative emotion, story moral, or narrative intent. We label these features perspectival when there is a distribution of possible answers that is also contextually dependent. Finally, we use consensus where there are clear central tendencies in a distribution of answers.
@@ 5 Benchmark Survey
The purpose of the survey is to adjudicate to what degree currently available LLM benchmarks satisfy the aspects of narrativity identified above. For our purposes, we define a benchmark to be an annotated dataset or testing technique made available for testing language model accuracy on some qualitative task, with the following requirements. The benchmark should:
@@
Following these guidelines, we surface 78 potential benchmarks published over the last twelve years. Filtering for papers that provide their code and data in public repositories reduces this count to 39 benchmarks. We then adjudicate how well each benchmark matches the respective evaluation attributes for each hierarchal attribute triple (e.g. local-discrete-deterministic), measuring the edit distance. From this value we assign alignment ratings with our taxonomy of “good,” (a distance of 0), “decent,” (1), “poor,” (2), and finally “bad” (3), resulting in the distribution seen in Table 1. Retaining all benchmarks with a quality rating of “poor” and above leaves us with a total 34 benchmarks which we present in Figure 3 and with a complete rating breakdown in Table 3. We now highlight various trends we observed when conducting our survey. These trends are both positive and negative.
@@
Of the 39 benchmarks surveyed, 27 rely on classification or multi-label prediction, while 12 evaluate open-ended text generation. Large language models are, at their core, generative systems. Evaluating their generated text therefore offers a direct window into their underlying biases and proficiencies—a method increasingly adopted in recent work (Lucy and Bamman, 2021; Hicke et al., 2025). Expanding such generative benchmarking approaches would enable the study of model responses in more holistic and perspectival ways.
@@
We identified nineteen benchmarks that address story-specific features in our taxonomy. Compared to benchmarks identified for narration (2), discourse (5), and situatedness (5), this suggests current benchmarking efforts overemphasize the story as a target. One of the principal aims of our theoretical framework is to foreground the complexity of narrative communication and the importance of narrative perspective, discourse structure, and social context for surfacing narrative meaning.
@@
While we identified many existing annotated datasets capturing event-related tasks (Baker et al., 1998; Sims et al., 2019; Tang et al., 2021; Wang et al., 2022b), none were expressly equipped for benchmarking as outlined in our requirements. These annotated datasets, while valuable, still need to be converted to the criteria outlined in the beginning of section 5 to be considered benchmarks.
@@
All but two of the benchmarks surveyed constructed their ground truth in a deterministic manner, meaning LLMs must respond with the correct answer from a set of possible responses to be considered correct. As we note in section 2, benchmark developers are now phasing this mode of judgment out in favour of allowing for a multiplicity of responses. This method is more accurate when considering ground truth is often based on annotations from multiple annotators, meaning ground truth begins as a distribution of responses. Non-deterministic responses likewise more proficiently capture the subjective quality of audience response, in that no two audience members will respond to a narrative in precisely the same way. The same expectation must be afforded to LLMs if their abilities are to be judged in a realistic manner.
@@
Early statistical NLP techniques were often built on lexical corpora attributing per-token scores for various linguistic phenomena. This enabled algorithms for assessing the phenomena over particular substrings, important for researchers interested in identifying unexpected failure modes over text input Lin (2004). The convenience of passing in an entire text to a LLM and receiving back a comprehensive answer has motivated benchmark developers to prefer holistic and global assessments over per-token discrete annotations. This has the benefit of allowing language models the space to respond as they naturally would to a query of a similar nature, but evaluating per-token responses can elicit a more nuanced understanding of how language models represent complex semantics. We argue both approaches are important for narrative understanding and encourage benchmark developers to continue to value token-level benchmarking.
@@
Four of the 39 benchmarks we shortlisted were multilingual, with the rest only implementing their test in English. This language gap leaves open the question of whether language models can properly represent narrative dimensions in non-English languages (and especially low-resource languages). We encourage benchmark developers to seek out platforms that provide human assessments from global perspectives and resources that offer narrative data in multiple languages.
@@
Narratives exist in multiple modalities. Narratological research in recent decades has developed theoretical frameworks for describing how narratives behave in music, imagery, paintings, and movies (Fludernik, 2010). Approximately 5% of our surveyed benchmarks were either multimodal or non-textual in their inputs, testing vision language models (VLMs) on their visual understanding abilities. Future benchmarks will want to combine these approaches to work towards a framework for describing narrative conceits in multiple mediums, especially as video-based social media platforms continue to grow in popularity.

### v3

@@ 1 Introduction
Narratives are omnipresent in daily life. They can be used to entertain, inform, persuade, and maintain shared beliefs in communities across generations. Designing large language models (LLMs) that can both understand and generate narrative communication is of utmost importance to building artificially intelligent systems useful for humans.
Figure 1: The twelve primary narrative features of the NarraBench taxonomy coloured according to the Big-4 narrative dimensions (story, narration, discourse, and situatedness) and shaded by how well existing benchmarks match as determined by our survey.
Using narratives to test long context understanding and models’ theory of mind has come into vogue of late (kimFANToMBenchmarkStresstesting2023; karpinskaOneThousandOne2024; chenMBENCHBenchmarkingTheory; wangNovelQABenchmarkingQuestion2025), and non-narrative-specific benchmarks can test for narrative understanding in unexpected ways: “Needle in a haystack” tests show models can discern what is relevant to the present contextual frame (kamradtNeedleHaystackPressure2025), and summarization tasks require models to integrate information across long contexts to identify important through-lines (linROUGEPackageAutomatic2004).
Several large-scale benchmarks contain a significant amount of narrative content in this way. For example, we estimate approximately 1 in 7 BIG-bench questions contains a story (srivastavaImitationGameQuantifying2023), 1 in 4 in HellaSwag (zellersHellaSwagCanMachine2019), and 1 in 3 in MMLU (hendrycksMeasuringMassiveMultitask2021). [^1] All datasets were classified on a per-document basis with the BERT-based classifier from antoniakWherePeopleTell2024. Correctly answering these questions indicates some level of narrative understanding.
Nevertheless, benchmarks that explicitly and systematically test for narrative understanding are rare and often lack theoretical consistency in their principal goals and methodological foundations. Existing benchmarks often deploy single-answer frameworks, overlooking the way narrative understanding is a fundamentally interpretive (i.e., perspectival, not deterministic) task. Benchmarks likewise tend to focus on story-level content, missing more complex dimensions of narrative communication (e.g. discourse structure and narrative perspective).
@@ 2 Related Work
Early computational work modelled narratives as stochastic processes (kahnFinitestateModelsPlot1973) or scripts known as Fillmorean frames (schankScriptsPlansGoals1975) containing chronological chains of events composed of entities and their actions (chambersUnsupervisedLearningNarrative2008; chambersUnsupervisedLearningNarrative2009; chambersDatabaseNarrativeSchemas2010; balasubramanianGeneratingCoherentEvent2013; reiterNLPbasedCrossdocumentApproach2014). Advancements in NLP led to increasingly sophisticated narrative extraction methods that made use of topic models, manually derived features, and associated rule mining (antoniakNarrativePathsNegotiation2019a; belyyScriptInductionAssociation2020; lyuGoalOrientedScriptConstruction2021; zhangHumanintheLoopSchemaInduction2023). This approach was extended to plot arc extraction, newly reinterpreted as sentiment over time (hoganAffectiveNarratologyEmotional2011; reaganEmotionalArcsStories2016; somasundaranEmotionArcsStudent2020; elkinsShapesStoriesSentiment2022; knight2024narrative), and social network analysis, where entities and their actions were rendered as graphs (roemmeleIdentifyingSensibleLexical2019; simsMeasuringInformationPropagation2020a; tangherlini2020automated). Further work has focused on higher-level understanding including narrativity detection antoniakWherePeopleTell2025; piper-bagga-2025-narradetect and narrative intent understanding zhuAreNLPModels2023. As LLMs are increasingly used to assess and generate narratives, now is an optimal time to codify the diverse array of tasks that fall under the heading of narrative understanding.
@@
Benchmarks are the principal infrastructure of empirical NLP, enabling standardized comparison and shared progress measures across models. Of particular interest has been benchmarks for evaluating long-context understanding (changBooookScoreSystematicExploration2024; karpinskaOneThousandOne2024; wangNovelQABenchmarkingQuestion2025; thaiLiteraryEvidenceRetrieval2025) and summarization performance (fabbriSummEvalReevaluatingSummarization2021; zhaoNarraSumLargeScaleDataset2022; subbiahSTORYSUMMEvaluatingFaithfulness2025). A smaller set of benchmarks focus on explicitly narrative phenomena—including dialogue (vishnubhotlaProjectDialogismNovel2022), plot arcs (chunSentimentArcsNovelMethod2021a), or genre classification (worshamGenreIdentificationCompositional). General benchmarks for narrative understanding remain scarce. Evaluation methods also remain limited; most rely on multiple-choice questions for efficiency — increasingly seen as unreliable estimates of model comprehension (sapSocialIQaCommonsense2019; wangMyAnswerFirstToken2024a; balepurArtifactsAbductionHow2024; alzahraniWhenBenchmarksAre2024; wangLLMsMayPerform2025). Recent alternatives such as LLM-as-a-Judge evaluation (liLLMsasJudgesComprehensiveSurvey2024; tanJudgeBenchBenchmarkEvaluating2025; wangImprovingLLMasaJudgeInference2025) and distributional or multi-annotator scoring (niMixevalDerivingWisdom2024; meisterBenchmarkingDistributionalAlignment2024) point toward richer, less rigid evaluation pipelines, but have not yet been applied to narrative domains.
@@ 3 Theoretical Foundations
Narrative theory has developed, over more than a century, increasingly sophisticated qualitative descriptions of narratives across different mediums and contextual settings.
@@ 3.1 The Elements of Narrative
Our base definition of narrative communication is drawn from piperNarrativeTheoryComputational2021. At its most elementary level, a narrative can be said to occur when all of the following criteria are met:
@@ A Someone
Narrative theory has additionally foregrounded the centrality of change or conflict as an elementary component of narrative communication herman2009basic, through concepts like “change of state” prince_narratology_2012, “canonicity/breach” bruner_narrative_1991, denouement freytag1895technique, or Aristotle’s foundational emphasis in the Poetics (c. 330 bce) on beginning, middle, and end structure. At a root level, narratives foreground temporal difference with respect to agent-centred events genetteNarrativeDiscourseEssay1980; sternberg1992telling; ricoeur2012time.
@@ 3.2 Higher-Order Narrative Dimensions
Figure 2: NarraBench’s primary theoretical foundations from genetteNarrativeDiscourseEssay1980 and herman2009basic.
To integrate these elements into a broader framework, we combine genetteNarrativeDiscourseEssay1980’s narrative triangle with herman2009basic’s notion of situatedness (Figure 2). The Russian formalists tomashevsky1965russian were the first to distinguish between story (“what happened”) and discourse (“how it was told”), to which genetteNarrativeDiscourseEssay1980 added a third dimension called narration, which refers to the perspectival dimensions of narrative communication produced by the narrator’s voice (“who speaks”).
Genette then introduced three further terms to capture the relationship between these dimensions, tense, mood, and voice. Genette extrapolates from the linguistic meanings of these terms to capture specific narratological features. These include aspects of time and the ordering of events (tense); the relationship between eventfulness and description (mood); and aspects related to perspective, such as point of view, dialogue, and focalization (voice). herman2009basic importantly adds a fourth dimension, situatedness, which captures the social dimensions of narrative (medium, social context, interactive dimensions). We combine these elementary and higher-level dimensions to structure our taxonomy described in the next section.
@@ 4 The NarraBench Taxonomy
Dimension Feature Aspect SMV
Story Agent name LDD
GHC
role GHP
attributes LDD
GHC
emotions LDP
GHP
motivation LDP
GPP
Social Net interaction LDD
connections GHD
relationship GHC
Event event LDD
GDC
schema GHC
causality GPP
Plot topic GHC
plot GHP
plotline GHC
moral GHP
obstacle GHP
conflict GHP
archetype GHC
structure plot arc GPC
Setting setting LDD
GHC
location LDD
GDD
Discourse Time duration LDD
GPD
GHD
order GPD
GHD
Revelation suspense GPP
curiosity GPP
surprise GPP
Narration Perspective point of view GDD
focalization LDD
dialogue LDD
Style allusion LDP
figurative LDP
imageability LHP
complexity LHP
evaluative LDP
Situatedness Paratext genre GHC
author GDD
date GDD
medium GDD
platform GDD
Motivation intent GHP
Table 1: The complete NarraBench taxonomy. We indicate SMV (Scale-Mode-Variance) with: L/G=local/global, D/H/P=discrete/holistic/progressive, D/C/P=deterministic/consensus/perspectival.
@@
The second level of our taxonomy focuses on evaluation criteria. These include textual scale (local, global, meso), mode (discrete, progressive, holistic judgments), and the expected variance of potential answers (deterministic, consensus, perspectival). We present the taxonomy in its entirety in Table 4, and a condensed version in Table 1.
@@ 4.1.1 STORY
Story captures all aspects of a narrative that relate to storyworld content. We pay attention to:
@@
Narratives foreground the lived experiences of agentic entities (i.e., characters) fludernik2002towards. Agent detection and inventorying character names are an essential first step. Higher-level synthesis of characters around their narrative roles and attributes is necessary for understanding the functional position of characters within narratives proppMorphologyFolktaleSecond1968. We also include theory-of-mind related tasks such as narrative emotion labelling and character motivation understanding. Interpreting agent intentionality and inner experience is a central component of narrative understanding zunshine2006we; mar2011neural; oatley2016fiction.
@@
Agents in a narrative exist in a particular social configuration. Here we divide social networks into three core aspects: interactions (local social events), connections (who does each character know), and relationships (the social types of connections, including employment, family, civil, romantic, and friendship-based connections) labatut2019extraction.
@@
Narratives are structured around agents who experience change over time. Such change arises through action, where an action and its corresponding effect together form a causal frame. While narratologists debate the precise boundaries of a frame within a story, it is broadly accepted that reconstructing the underlying chain of actions and their causal relations is central to recovering the narrative’s event structure or schemas schankScriptsPlansGoals1975; chambersUnsupervisedLearningNarrative2009.
@@
According to Kukkonen2014Plot, “The term plot designates the ways in which the events and characters’ actions in a story are arranged and how this arrangement in turn facilitates identification of their motivations and consequences.” In a technical sense, plot sits between story and discourse, but for our purposes we consider it as part of the story dimension. Plot can thus be subdivided into topics (core thematic concerns), plot lines (sub-plots), key structural forces like narrative obstacles or conflicts, central value lessons (i.e., morals), as well as larger schematic structures like archetypes.
@@
Kukkonen2014Plot makes a core distinction between plot as a function of “global structure” (our primary definition above) and “progressive structuration,” that is, aspects of plot that depend on the sequencing of events for their meaning. We reserve these progressive aspects for the discourse dimension, with the exception of plot arcs, which capture structural elements related to events such as fortune reaganEmotionalArcsStories2016; elkinsShapesStoriesSentiment2022; mcadams2001bad, reversal knight2024narrative, and denouement freytag1895technique. Understanding the reliance or inability of LLMs to produce certain plot arc structures is important for shaping reader expectations.
@@
The spatial distribution of perception and movement within narratives is a key aspect of defining audiences’ cognitive orientation. Where narrative theory has introduced concepts such as the “chronotope” bakhtin2010dialogic or “possible worlds” ryan2016narrating, we use the more integrative term of setting to capture both the holistic understanding of narrative space and the particular location of a given narrative moment.
@@ 4.1.2 NARRATION
Narratives require narrators. According to genetteNarrativeDiscourseEssay1980, “who speaks” plays a central role in governing the selection and communication of narrative content, which will in turn impact how audiences interpret the narrative.
@@
We use style to capture rhetorical and intertextual aspects of narratives that bear on their meaning. Allusion refers to inter-textual references kristeva1980desire; genette1997palimpsests, while figurative language and imageability are concerned with the symbolic and/or perceptual dimensions of narrative communication. We also include complexity as a measure of reading level and social capital and evaluative language as a way of capturing narrative meta-awareness labov1997narrative.
@@ 4.1.3 DISCOURSE
Narratives take place over time. But time is not necessarily represented in a chronological manner, nor is time represented in an equitable manner. Narrators may decide to dedicate greater fidelity to certain events, or reflect back on prior events or foreshadow future events. Following genetteNarrativeDiscourseEssay1980, we focus on both duration and order with respect to time.
@@
How information is revealed to the reader over the course of narrative time fundamentally shapes readers’ affective state brewer1982stories. Here we focus on the classic tripartite scheme of reader suspense, curiosity, and surprise to capture the assessment of narrative revelation over time.
@@ 4.2 Evaluation Criteria
This dimension captures assumptions about interpretive freedom. Some features may require single, correct answers, such as a character name. We label these deterministic. Other features are more open-ended, such as narrative emotion or narrative intent. We label these features perspectival when there is a distribution of possible answers that is also contextually dependent. We finally use consensus where there are clear central tendencies in a distribution of answers.
@@ 5 Benchmark Survey
The purpose of the survey is to adjudicate to what degree currently available LLM benchmarks satisfy the aspects of narrativity identified above. For our purposes, we define a benchmark to be an annotated dataset or testing technique made available for testing language model accuracy on some qualitative task. The benchmark should:
@@
Following these guidelines, we surface 78 potential benchmarks published over the last twelve years via relevant keyword searches on Google Scholar. Filtering for open data reduces this count to 39 benchmarks. We then adjudicate how well each benchmark matches the respective evaluation attributes for each hierarchal attribute triple (e.g. local-discrete-deterministic), measuring the edit distance. From this value we assign alignment ratings with our taxonomy of “good,” (a distance of 0), “decent,” (1), “poor,” (2), and finally “bad” (3), resulting in the distribution seen in Table 2. Retaining all benchmarks with a quality rating of “poor” and above leaves us with 34 benchmarks which we present in Figure 3 and with a complete rating breakdown in Table 5. We now highlight trends observed when conducting our survey.
@@
Of the 39 benchmarks surveyed, 27 rely on classification or multi-label prediction, while 12 evaluate open-ended text generation. Causal LLMs are generative models. Evaluating their generated text therefore offers a direct window into their underlying biases and proficiencies, as has been increasingly adopted (lucyGenderRepresentationBias2021; hickeZeroBodyProblem2025). Expanding such generative benchmarking approaches would enable the study of model responses in more holistic and perspectival ways.
Attributes Emotional Connections Plotline Order Order Order
Model austenalike culemo phantomwiki storysumm tot tram traveler Mean
Gemma-3 12B 3.2 41.0 17.8 40.6 3.8 67.0 54.0 32.5
GPT-OSS 20B 0.0 100.0 2.6 62.5 1.2 0.0 19.0 26.5
Llama-3 8B 2.1 28.5 23.0 55.2 31.8 59.5 61.0 37.3
Table 3: Results across all well-conforming benchmarks and models. Accuracy is reported as percentages.
@@
We identified 19 benchmarks that address our story-specific features. Compared to benchmarks identified for narration (2), discourse (5), and situatedness (5), this suggests current benchmarking efforts overemphasize the story as a target. One of the principal aims of our theoretical framework is to foreground the complexity of narrative communication and the importance of narrative perspective, discourse structure, and social context for surfacing narrative meaning.
@@
While we identified many existing annotated datasets capturing event-related tasks (bakerBerkeleyFrameNetProject1998; simsLiteraryEventDetection2019; tangDiscourseNarrativeKnowledge2021; wangUncoveringSurprisingEvent2022), none were expressly equipped for benchmarking as outlined in our requirements. These annotated datasets, while valuable, still need to be converted to the criteria outlined in the beginning of section 5 to be considered benchmarks.
@@
All but two of the benchmarks constructed their ground truth in a deterministic manner, meaning LLMs must must correctly select an answer from a set of possible responses. As we note in section 2, benchmark developers are now phasing this mode of judgment out in favour of allowing for a multiplicity of responses. This method is more accurate when considering ground truth is often based on annotations from multiple annotators, meaning ground truth begins as a distribution of responses. Non-deterministic responses likewise more proficiently capture the subjective quality of audience response, in that no two audience members will respond to a narrative in precisely the same way. The same expectation must be afforded to LLMs if their abilities are to be fairly judged (plank-2022-problem; pichler-etal-2025-evaluating).
@@
Early statistical NLP techniques were often built on lexical corpora attributing per-token scores for various linguistic phenomena. This enabled algorithms for assessing the phenomena over particular substrings, important for researchers interested in identifying unexpected failure modes over text input linROUGEPackageAutomatic2004. The convenience of passing in an entire text to a LLM and receiving back whole answers has motivated developers to prefer holistic/global assessments over discrete responses. This has the benefit of allowing language models the space to respond as they naturally would to a query of a similar nature, but evaluating per-token responses can elicit a more nuanced understanding of how language models represent complex semantics. We argue both approaches are important for narrative understanding and encourage benchmark developers to continue to value token-level benchmarking.
@@
Four of the 39 benchmarks we shortlisted were multilingual, with the rest implementing their test in English. This gap leaves open the question of whether language models can properly represent narrative dimensions in non-English languages (and especially low-resource languages). We encourage benchmark developers to seek platforms that provide human assessments from global perspectives alongside multilingual narrative resources.
@@
Narratives exist in multiple modalities. Narratological research in recent decades has developed theoretical frameworks for describing how narratives behave in music, imagery, paintings, and movies (fludernikPostclassicalNarratologyApproaches2010). Approximately 5% of our surveyed benchmarks were either multimodal or non-textual in their inputs, testing vision language models (VLMs) on their visual understanding abilities. Future benchmarks will want to combine these approaches to work towards a framework for describing multimodal narrative concepts.
6 Testing Harness
NarraBench is a novel attempt at reconfiguring existing benchmarks to fulfill theoretical goals, and as such can be implemented as an expandable collection of relevant benchmarks. To guide future implementations we have constructed a testing harness containing the most suitable benchmarks identified in our study, showcasing sample scores from a series of popular open-weight LLMs in Table 3.222We make our harness available here. Our results suggest non-uniform performance across models and specific narrative aspects, a result not immediately obvious when making use of a specific benchmark as an indicator for story understanding. We encourage all interested to identify or develop further benchmarks in areas not yet covered as per our initial survey. Distributing a stable implementation is one important way NarraBench will remain an evolving and relevant resource moving into the future.